OpenStudy (anonymous):

Which of the following are solutions to (2tanx)/(1 - tan^2x) = √3 ? Check all that apply.

OpenStudy (anonymous):

A. x = 5π/6 + nπ B. x = π/3 + nπ C. x = π/6 + nπ D. x = 2π/3 + nπ

OpenStudy (aravindg):

(2tanx)/(1 - tan^2x) =tan 2x

OpenStudy (aravindg):

Can you do it now

OpenStudy (anonymous):

Is that not an identity?

OpenStudy (aravindg):

Yes.

OpenStudy (anonymous):

But it can't be all of them, right...

OpenStudy (aravindg):

It cant be all of them. tan 2x=root(3) Solve for x

OpenStudy (anonymous):

From what I've got, its only C.

OpenStudy (aravindg):

You are right

OpenStudy (anonymous):

C and D